# zip2zip: Inference-Time Adaptive Vocabularies for Language Models via Token Compression

### v1

@@ 6 Conclusion
We introduced zip2zip, a framework for inference-time vocabulary adaptation with LLMs. By integrating LZW-based token compression with a dynamic hypertoken embedding mechanism, zip2zip enables substantial reductions in sequence length and decoding steps, leading to improved inference efficiency with minimal architectural modifications. Our experiments demonstrate that zip2zip maintains strong performance across a range of tasks while achieving significant gains in inference efficiency. These findings highlight the promise of integrating dynamic tokenization into LLMs, opening up new directions for research in LLM efficiency.
Appendix A Ablation Studies
Definition A.1 (Compression Rate).
@@ Appendix A Ablation Studies
Compression $\text{Compression Rate}=\frac{N_{\text{comp}}}{N_{\text{orig}}}\times 100\%.$
@@ A.1 LZW Maximum Merge Size
A.2 Hyper-encoder architecture
@@
Appendix B FLOPs Estimation for zip2zip
@@
Appendix C Additional Results
Machine Translation
@@ Appendix C Additional Results
Table 10: Standard deviations (bootstrapped) for machine translation scores across WMT benchmarks.
@@
Phi-3-4B +Raw ±2.1 ±1.4 ±1.7 ±1.9 ±1.6 ±1.8 ±1.5 ±1.3 ±2.4
+Finetune ±2.2 ±1.6 ±1.8 ±1.8 ±1.4 ±1.7 ±1.4 ±1.5 ±2.3
+zip2zip ±1.9 ±1.5 ±2.0 ±1.7 ±1.6 ±1.9 ±1.6 ±1.4 ±2.5
Phi-3-14B +Raw ±2.0 ±1.4 ±1.9 ±2.0 ±1.5 ±1.7 ±1.5 ±1.4 ±2.2
+Finetune ±2.2 ±1.4 ±1.9 ±2.0 ±1.3 ±1.9 ±1.4 ±1.3 ±2.9
+zip2zip ±2.1 ±1.5 ±1.8 ±2.1 ±1.6 ±1.8 ±1.5 ±1.3 ±2.6
Appendix D Technical Details
Model and Training Configuration
@@ Appendix D Technical Details
System and Libraries
@@
Appendix E Data Mixture
@@
Appendix F Token Stream Visualization
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We have clearly stated the claims, assumption and contributions we have made in the abstract and introduction.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: We have included a section called Discussion and Limitations in the main text.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: While we don’t theory results. We have tried to formalize our method with formulas but it’s not really theoretical contribution. And we have tried to make it clear with empirical results.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: We have described how our methods works in details and we also provide code and specify dataset and experimental details in the supplementary material as well as on this anonymous repository.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: We provide code and data, as well as detailed experimental settings, in the supplementary material, as well as on this anonymous repository.
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: It can be found in the code included in the supplemental material; also available on this anonymous repository. We also include the settings of key hyperparameters in Appendix D.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: For brevity, in the main text we report experimental results in tables, and we leave the corresponding standard deviations in Table 10. For the two-shot accuracy, we report the standard deviations in the title of Table 4, which are approximately always 0.02.
  29. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  30. 8.
    Experiments compute resources
  31. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  32. Answer: [Yes]
  33. Justification: We provide detailed information on the computer resources in Appendix D.
  34. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  35. 9.
    Code of ethics
  36. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  37. Answer: [Yes]
  38. Justification: We have reviewed the NeurIPS Code of Ethics, and we confirm that the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
  39. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  40. 10.
    Broader impacts
  41. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  42. Answer: [N/A]
  43. Justification: This work does not involve any societal impacts.
  44. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  45. 11.
    Safeguards
  46. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  47. Answer: [N/A]
  48. Justification: The paper does not involve any high-risk components such as pretrained generative models or web-scraped data.
  49. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  50. 12.
    Licenses for existing assets
  51. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  52. Answer: [Yes]
  53. Justification: We have cited the original paper that produced the code package, dataset, and models, and included the url links when applicable.
  54. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  55. 13.
    New assets
  56. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  57. Answer: [Yes]
  58. Justification: We have provided our code package and detailed settings, as well as license, in this anonymous link.
  59. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  60. 14.
    Crowdsourcing and research with human subjects
  61. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  62. Answer: [N/A]
  63. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  64. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  65. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  66. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  67. Answer: [N/A]
  68. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  69. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  70. 16.
    Declaration of LLM usage
  71. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  72. Answer: [N/A]
  73. Justification: The core method development in this paper does not involve any LLM as any important, original, or non-standard components.
  74. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@ 6 Conclusion
We presented zip2zip, an approach that brings inference-time tokenization to large language models through online token compression. By combining LZW-based sequence compression with dynamic hypertoken embeddings, zip2zip enables compact, adaptive tokenization with lightweight uptraining and little architectural changes. Across multiple domains and languages, it achieves substantial reductions in sequence length and decoding cost while maintaining strong task performance. These results demonstrate that online token compression can serve as a practical path toward dynamic tokenization, pointing to new directions for efficient and adaptable LLM inference.
Acknowledgements
We would like to thank Miao Xiong for proof-reading the paper and providing valuable feedback. We are also grateful to Emre Kıcıman, Jason Eisner, Barun Patra, Ana-Maria Indreias, Xiuying Wei, Julian Minder, Tiago Pimental, Luca Beurer-Kellner, and Aleksei Kudrinskii for their helpful input and insightful discussions throughout this project. Special thanks to Zheng Zhou for providing technical support with the computing infrastructure.
West’s lab is partly supported by grants from Swiss National Science Foundation (TMSGI2_211379 and Grant 200364), Swiss Data Science Center (P22_08), H2020 (952215), Microsoft Swiss Joint Research Center, and Google, and by generous gifts from Facebook, Google, and Microsoft.
Appendix A More Related Work
Wang et al. [2025], Liu et al. [2024a] conducted studies on how to effectively expand the vocabulary by better selecting the subset of tokens to add. Multi-Word Tokenizer [Gee et al., 2023] and SuperBPE [Liu et al., 2025] demonstrated that allowing forming tokens beyond word boundaries in BPE vocab learning helps to achieve more compact tokenization and even improves model performance. Content-Adaptive Tokenizer (CAT) by  [Shen et al., 2025] introduces a dynamic image tokenization approach that allocates tokens based on content complexity, achieving improved reconstruction quality and efficiency compared to fixed-size tokenization methods. [Lotz et al., 2025] propose a pixel-level fallback encoder that bypasses subword vocabulary limitations by rendering text as images, enabling vocabulary-free representations that improve multilingual performance and efficiency in pretrained language models. The FlexiTokens [Owodunni et al., 2025] introduces learnable, byte-level tokenizers that dynamically adapt token boundaries to new domains and languages, reducing over-fragmentation.
Appendix B Discussions on Merge Size
B.1 An Upper Bound on Merge Size
Proposition B.1.
Let $T$ be an input sequence of length $N$ over a finite alphabet. LZW compression algorithm merges substrings by identifying and replacing the most frequent substrings with new symbols, iteratively. Then, the size $M$ of the largest merged unit (i.e., the longest substring created via merging) is bounded above by $O(\sqrt{N})$.
Proof.
Assume that the largest merged unit has size $M$. This implies that there exists at least one merge at level $M$ involving a substring of length $M$. Furthermore, due to the nature of merge-based algorithms, any merged unit of size $k$ must be composed of previously merged units of smaller sizes (e.g., from sizes k1k-1 and $1$, or similar). Hence, in order to construct a merged unit of size $M$, the algorithm must have previously created all merged units of sizes $1$ through M1M-1.
Thus, the existence of a merged unit of size $M$ implies the existence of merged units of every size $k$ such that $1\leq k\leq M$. Each such unit must occur at least once in the input sequence in order to be merged.
Therefore, the total number of characters in $T$ must be at least the sum of the lengths of all merged units from size $1$ to $M$, i.e.,
$N\geq\sum_{k=1}^{M}k=\frac{M(M+1)}{2}.$
This implies:
$M=O(\sqrt{N}).$
Thus, the length $M$ of the largest merged unit is bounded above by $O(\sqrt{N})$. ∎
B.2 Relation Between Merge Size and Compression Rate
Definition B.1 (Compression Rate).
@@ B.2 Relation Between Merge Size and Compression Rate
$\text{Compression Rate}=\frac{N_{\text{comp}}}{N_{\text{orig}}}\times 100\%.$
@@
Appendix C Discussions on Hyper-Encoder Architecture
C.1 Hyper-encoder architecture
@@
Appendix D Discussions on Compression
Table 8: Token statistics for Code Generation, Biomedical, and French QA domains.
Stats Code Biomedical French
Original Seq Len 344 322 399
Zip2Zip Seq Len 265 270 367
Num Hypertoken 44 35 26
Hypertoken Ratio 0.166 0.130 0.071
Compression Rate 0.770 0.839 0.920
Table 8 shows detailed token statistics on illustrative examples across three domains in Figure 4, highlighting zip2zip’s ability to reduce sequence length and introduce reusable hypertokens with domain-specific efficiency.
Table 9: Sequence length reduction (%) across domains, inferred from the inverse of token efficiency gains in Table 2.
Tokenizer Code Math Chat Multilingual Web
Llama-3-128K 34.9% 32.5% 20.3% 19.1% 14.8%
Qwen-2-150K 35.5% 37.8% 20.3% 19.6% 15.4%
Phi-4-200K 34.9% 34.1% 19.4% 16.4% 13.0%
Gemma-3-256K 41.1% 37.8% 21.9% 18.5% 16.7%
Table 9 reports the estimated sequence length reduction across domains, showing that zip2zip consistently shortens token sequences by 13–41%, with the strongest gains observed in structured domains like code and math.
Appendix E FLOPs Estimation for zip2zip
@@
Appendix F Discussions on Tokenizer
Figure 7 compares the tokenization and detokenization latencies across different tokenizer configurations. The Base Tokenizer corresponds to the standard BPE tokenizer implemented by the Hugging Face tokenizers library. The Rust LZW Tokenizer represents the end-to-end latency when LZW compression and decompression are applied on top of the BPE tokenization. As shown, this configuration introduces only a small additional latency in the tokenization process while leaving the detokenization latency virtually unchanged. The Python LZW Tokenizer, in contrast, exhibits significantly higher latency due to Python’s runtime overhead. Overall, the results indicate that most of the observed latency arises from the BPE segmentation process itself rather than the LZW compression, suggesting that efficient implementations of compression add minimal overhead to tokenization workflows.
Figure 7: zip2zip tokenizer latency (ms) vs. HF tokenizer.
Appendix G Entropy Invariance under Lossless Transformations
Theorem G.1 (Entropy Invariance under Lossless Compression).
Let $g:\mathcal{X}^{\ast}\to\mathcal{Z}^{\ast}$ be a bijection onto its image, and let $P_{Z}$ and $p_{\gamma}$ denote the push-forward distributions of $P_{X}$ and $p_{\theta}$, respectively:
$P_{Z}(z)=P_{X}(g^{-1}(z)),\qquad p_{\gamma}(z)=p_{\theta}(g^{-1}(z)).$
Then the total entropy and cross-entropy are invariant under $g$:
$H(P_{Z})=H(P_{X}),\qquad H(P_{Z},p_{\gamma})=H(P_{X},p_{\theta}).$
$\mathcal{X}^{\ast}$$\mathcal{Z}^{\ast}$$x_{1}$$x_{2}$$x_{3}$$x_{4}$$z_{1}$$z_{2}$$z_{3}$$z_{4}$$g:\mathcal{X}^{\ast}\to\mathcal{Z}^{\ast}$
Figure 8: Lossless compression mapping (bijective) $g$ (e.g., LZW) from original sequences $\mathcal{X}^{\ast}$ to compressed sequences $\mathcal{Z}^{\ast}$.
Proof.
By definition of the push-forward distribution,
$H(P_{Z})$ $=-\sum_{z\in\mathcal{Z}^{\ast}}P_{Z}(z)\log P_{Z}(z)$
$=-\sum_{z\in\mathcal{Z}^{\ast}}P_{X}(g^{-1}(z))\log P_{X}(g^{-1}(z))$
$=-\sum_{g(x^{\prime})\in\mathcal{Z}^{\ast}}P_{X}(x^{\prime})\log P_{X}(x^{\prime}).$
Since $g$ is a bijection, we can replace the summation by $x^{\prime}\in\mathcal{X}^{\ast}$, which yields
$H(P_{Z})=-\sum_{x\in\mathcal{X}^{\ast}}P_{X}(x)\log P_{X}(x)=H(P_{X}).$
An identical argument applies to the cross-entropy by considering $P_{X}(x)\log p_{\theta}(x)$ as a single function, since the invariance property does not depend on the specific form of the function being summed:
$H(P_{Z},p_{\gamma})$ $=-\sum_{z}P_{Z}(z)\log p_{\gamma}(z)$
$=-\sum_{x}P_{X}(x)\log p_{\theta}(x)=H(P_{X},p_{\theta}).$
Hence, both entropy and cross-entropy remain invariant under any lossless (bijective) transformation $g$. ∎
Appendix H Additional Results
H.1 Machine Translation
@@
Table 11: Machine translation performance on WMT benchmarks (BLEU↑, CHRF↑, TER↓) with standard deviations (±) from bootstrapped estimates. Scores are averaged across both directions.
@@
Phi-3.5-4B Base 33.6±2.1 58.3±1.4 53.0±1.7 39.2±1.9 63.2±1.6 47.9±1.8 17.7±1.5 45.5±1.3 73.4±2.4
Cont. pretrain 36.5±2.2 61.0±1.6 51.5±1.8 42.3±1.8 65.4±1.4 44.9±1.7 16.7±1.4 45.8±1.5 79.7±2.3
zip2zip 34.1±1.9 59.4±1.5 54.5±2.0 39.7±1.7 64.5±1.6 48.0±1.9 14.3±1.6 44.2±1.4 93.5±2.5
Phi-3-14B Base 39.1±2.0 62.6±1.4 49.3±1.9 43.1±2.0 65.6±1.5 44.1±1.7 21.3±1.5 51.0±1.4 70.5±2.2
Cont. pretrain 38.9±2.2 63.2±1.4 48.8±1.9 48.4±2.0 70.1±1.3 39.8±1.9 21.8±1.4 52.0±1.3 68.3±2.9
zip2zip 36.4±2.1 62.8±1.5 51.2±1.8 44.8±2.1 68.1±1.6 42.9±1.8 19.5±1.5 50.1±1.3 72.9±2.6
H.2 Multilingual QA Tasks
We evaluate Phi-3.5–4B on multilingual downstream tasks beyond translation, including TruthfulQA-2, HellaSwag, and Winograd across five languages (French, Spanish, Russian, Chinese, and Arabic). As shown in Table 12, the base model demonstrates strong cross-lingual generalization. Continued pretraining in the vanilla setting slightly reduces performance, likely due to domain drift, while the zip2zip variant performs similarly. Overall, the results indicate that the proposed multilingual adaptation strategy maintains competitive performance across diverse evaluation benchmarks.
Table 12: Evaluation on Multilingual Tasks in addition to Translation. We report results on TruthfulQA-2, HellaSwag, and Winograd across 5 languages (FR, ES, RU, ZH, AR) using lm-evaluation-harness.
Model Method TruthfulQA-2 HellaSwag Winograd
FR ES RU ZH AR FR ES RU ZH AR FR ES RU ZH AR
Phi-3.5–4B Base 0.47 0.51 0.46 0.41 0.42 0.61 0.66 0.49 NA 0.44 0.70 NA 0.74 0.74 NA
Cont. Pretrain. Vanilla 0.42 0.47 0.46 0.46 0.43 0.54 0.57 0.37 NA 0.27 0.76 NA 0.70 0.61 NA
Cont. Pretrain. zip2zip 0.41 0.46 0.47 0.43 0.40 0.55 0.55 0.36 NA 0.27 0.77 NA 0.73 0.64 NA
Appendix I Technical Details
I.1 Model and Training Configuration
@@ I.2 LoRA Configuration
I.3 Loss Weighting Coefficient
Since both the language modeling loss and the auxiliary reconstruction loss are formulated as cross-entropy objectives over tokens, they are naturally on a comparable scale. For a sequence of $N$ base tokens, the number of hypertokens used in the auto-reconstruction loss is upper-bounded by $N$, while each hypertoken corresponds to at most $M$ base tokens in the reconstruction target. The loss weighting coefficient $\lambda$ primarily serves to fine-tune the relative importance of the auxiliary objective rather than to correct for scale mismatch. We set $\lambda=0.1$, which yielded stable training. We did not perform extensive exploration over the value of $\lambda$, which we leave as an interesting direction for future work.
I.4 System and Libraries
@@ I.6 Evaluation
I.7 Loss Ratio
Appendix J Data Mixture
@@ I.7 Loss Ratio
Appendix K Token Stream Visualization